### v3

# Is Causality Necessary for Efficient Portfolios? A Computational Perspective on Predictive Validity and Model Misspecification

Abstract
A recent line of research has argued that causal factor models are necessary for portfolio optimization, claiming that structurally misspecified models inevitably produce inverted signals and nonviable frontiers. This paper challenges that view. We show, through theoretical analysis, simulation counterexamples, and empirical validation, that predictive models can remain operationally valid even when structurally incorrect. Our contributions are fourfold. First, we distinguish between directional agreement, ranking, and calibration, proving that sign alignment alone does not ensure efficiency when signals are mis-scaled. Second, we establish that structurally misspecified signals can still yield convex and viable efficient frontiers provided they maintain directional alignment with true returns. Third, we derive and empirically confirm a quantitative scaling law that shows how Sharpe ratios contract smoothly with declining alignment, thereby clarifying the role of calibration within the efficient set. Fourth, we validate these results on real financial data, demonstrating that predictive signals, despite structural imperfections, can support coherent frontiers. These findings refine the debate on causality in portfolio modeling. While causal inference remains valuable for interpretability and risk attribution, it is not a prerequisite for optimization efficiency. Ultimately, what matters is the directional fidelity and calibration of predictive signals in relation to their intended use in robust portfolio construction.
Keywords: Predictive modeling; Portfolio optimization; Causal inference; Misspecification; Mean-variance frontier; Signal calibration.
1 Introduction
Recent contributions in empirical finance have argued that portfolio optimization requires causal factor models, contending that structurally misspecified predictors, those that omit relevant variables or misrepresent functional relationships, necessarily generate distorted signals and collapse efficient frontiers (López de Prado and Zoonekynd,, 2024; López de Prado et al.,, 2024, 2025). Within this perspective, causal identifiability is presented as a prerequisite for viability: without it, signal inversion and optimization failure are presumed unavoidable.
This paper challenges the universality of that claim. Through theoretical analysis, controlled simulations, and empirical evidence, we demonstrate that predictive models can remain operationally effective for portfolio construction even when structurally misspecified. In particular, we show that signals derived from such models often preserve directional alignment with the true return vector, which is sufficient to sustain convex and non-degenerate efficient frontiers under classical mean–variance optimization (Almgren and Chriss,, 2005). Inefficiency, in practice, arises more frequently from miscalibration of signal magnitudes rather than from systematic sign inversion.
The contributions of this paper are fourfold. First, we formalize the distinction between directional agreement, ranking, and calibration, proving that directional correctness is necessary but not sufficient for efficiency. Second, we establish geometric results showing that structurally misspecified but directionally aligned signals preserve frontier convexity and optimization validity. Third, we extend the portfolios-from-sorts framework of Almgren and Chriss, (2005) by deriving a scaling law that links Sharpe ratio performance to the cosine alignment between true and surrogate signals. This refinement clarifies that correct ranking ensures membership in the efficient set, while calibration governs quantitative efficiency within it. Fourth, we validate these results using both synthetic data and S&P 500 equity returns, showing that predictive signals lacking causal identifiability can nonetheless support coherent and actionable frontiers.
These findings contribute to the broader debate in finance and econometrics concerning the relative roles of causality and prediction. As emphasized by Shmueli, (2010), explanatory and predictive models serve distinct purposes and should be evaluated against different criteria. Recent critiques have underscored the epistemic limitations of causal reductionism in financial systems: because markets are adaptive and reflexive, unidirectional causal inference may be unstable or conceptually ill-suited (Polakow et al.,, 2024). From this vantage, calibration and directional validity constitute more pragmatic benchmarks for assessing model utility in portfolio design than structural fidelity per se. Related work on model risk and empirical fragility further illustrates the dangers of overstating causal claims in environments characterized by adaptation and self-reference (Harvey and Liu,, 2014; Bailey et al.,, 2017).
We do not dispute the value of causal inference, which remains indispensable for structural attribution, regime monitoring, and stress testing (Wilcox and Gebbie,, 2014; Lo,, 2017). Rather, we caution against overstating its necessity for optimization. What ultimately matters for decision-making is not causal identifiability in the strict sense, but the directional fidelity and calibration of predictive signals relative to their intended application in portfolio construction. In this respect, predictive modeling provides a robust and operationally effective alternative, and often a complement, to causal-first approaches.
The remainder of the paper is structured as follows. Section 2 reviews related work on causal and predictive modeling in finance. Section 3 develops the theoretical framework, establishing results on directionality, ranking, and calibration under misspecification. Section 4 presents empirical validation through simulations and market data. Section 5 examines implications for portfolio theory and modeling practice. Section 6 concludes.
2 Literature Review: Predictive Validity vs. Causal Necessity in Portfolio Modeling
A growing strand of recent literature in asset pricing has emphasized the role of causal factor models as a prerequisite for portfolio efficiency. According to this view, models that omit causal structure, whether due to unobserved confounders, misspecified functional forms, or reliance on purely associational data, are expected to suffer from systematic directional failure, yielding distorted alphas and invalid optimization outcomes (López de Prado et al.,, 2024; López de Prado and Zoonekynd,, 2024; López de Prado et al.,, 2025).
Yet this causal necessity argument rests on assumptions that often prove restrictive in practice. In particular, it underestimates the operational value of predictive modeling, especially in noisy, high-dimensional financial environments where out-of-sample generalization may matter more than structural interpretability. The methodological distinction is well established. Shmueli, (2010) highlight the contrast between explanatory models, which seek to uncover causal relationships, and predictive models, which prioritize forecasting accuracy. Similarly, Breiman, (2001) describes the divide between the “data modeling” culture oriented toward interpretability and the “algorithmic modeling” culture that emphasizes predictive performance.
Finance magnifies this tension. Unlike controlled experimental sciences, financial markets rarely provide stable instruments, suffer from pervasive nonstationarity, and exhibit reflexive dynamics in which causal graphs may be both unstable and unobservable. In such settings, predictive validity frequently dominates causal alignment as the more pragmatic standard of model utility. Moreover, portfolio optimization is notoriously sensitive to estimation error. Seminal work by Best and Grauer, (1991), Chopra and Ziemba, (1993), and Britten-Jones, (1999) demonstrates that even small errors in expected return estimates can induce large swings in portfolio weights. To mitigate this fragility, the literature has developed a series of robustification strategies, including the Black–Litterman model (Black and Litterman,, 1992), constrained optimization approaches (Jagannathan and Ma,, 2003), and shrinkage-based estimators. While Michaud, (1989) introduced a resampling approach where shrinkage was only a byproduct, the more direct lineage of shrinkage estimators originates with the James–Stein estimator (James and Stein,, 1961) and modern covariance shrinkage techniques such as those of Ledoit and Wolf, (2004). Further, DeMiguel et al., (2009) show that simple heuristics, such as equal-weighted portfolios, can outperform optimized allocations when estimation error is severe. Collectively, these results underscore the practical limitations of strictly structural models.
Against this backdrop, predictive-first approaches have gained momentum. Gu et al., (2020) demonstrate that machine learning methods, despite their lack of causal interpretability, deliver superior forecasts and robust portfolio allocations. Similarly, Haugen and Baker, (1996) and Harvey et al., (2016) provide evidence that return-predictive signals can remain operationally effective even without causal grounding. In parallel, researchers have advanced new dependence measures designed to overcome the shortcomings of correlation-based models without appealing to causal structure. For example, Opdyke, (2021) proposes positive definite dependence measures that enable richer associational modeling and improved stress-testing, while Viole and Nawrocki, (2016) develops partial-moment dependence metrics that further broaden the class of non-causal approaches to portfolio design. Together, these contributions demonstrate that non-causal dependence modeling constitutes a viable and expanding alternative to fully specified structural methods.
While causal inference retains value for specific tasks such as risk decomposition or regime identification, the broader evidence indicates that it is not a universal prerequisite for portfolio efficiency. The decisive question is whether a model generates signals that are directionally accurate, properly scaled, and operationally useful for allocation. Under this perspective, predictive validity is often sufficient, and, in many practical settings, more effective, than strict causal specification. The next sections formalize this claim by analyzing the conditions under which predictive signals remain viable despite structural misspecification, and by quantifying how calibration, rather than causal structure, governs realized efficiency.
3 Methods
3.1 Preliminary Work: Analyzing Misspecification in Linear Factor Models
This section analyzes the mathematical consequences of model misspecification on portfolio efficiency, focusing on linear factor models with omitted confounders. We examine how such structural misspecification introduces bias in estimated factor loadings and how this bias propagates to portfolio weights, even under valid optimization procedures.
3.1.1 Bias from Omitting a Confounder
Assume asset returns follow a true structural factor model:
$X=B_{\text{true}}F+\varepsilon$
where:
$B_{\text{true}}=\begin{bmatrix}\gamma_{1}&\beta_{1}\\ \gamma_{2}&\beta_{2}\end{bmatrix},\quad F=\begin{bmatrix}Z\\ F_{2}\end{bmatrix}$
Here, $Z$ is an unobserved confounder that affects both the returns and the observed factor $F_{2}$. Suppose:
$F_{2}=\delta Z+\eta,\quad\eta\sim\mathcal{N}(0,1-\delta^{2}),\quad Z\sim\mathcal{N}(0,1)$
with $Z$ and $\eta$ independent, so that111(López de Prado et al.,, 2024) incorrectly report $\mathrm{Var}(F_{2})=1+\delta^{2}$, which would be the case only if $\eta\sim\mathcal{N}(0,1)$. With their definition $\eta\sim\mathcal{N}(0,1-\delta^{2})$, the correct result is $\mathrm{Var}(F_{2})=1$. This difference affects the exact expression for the bias in $\hat{\beta}_{n}$: under their assumption $\hat{\beta}_{n}=(\beta_{n}+\gamma_{n}\delta)/(1+\delta^{2})$, whereas under the correct normalization $\hat{\beta}_{n}=\beta_{n}+\gamma_{n}\delta$. In both cases, however, omitting $Z$ induces a $\delta$-dependent bias, so the qualitative conclusion remains unchanged.:
$\text{Var}(F_{2})=\delta^{2}\cdot\text{Var}(Z)+\text{Var}(\eta)=\delta^{2}+(1-\delta^{2})=1,\quad\text{Cov}(Z,F_{2})=\delta.$
Now, regress $X_{n}$ on $F_{2}$ alone. The estimated loading becomes:
$\hat{\beta}_{n}=\frac{\text{Cov}(X_{n},F_{2})}{\text{Var}(F_{2})}=\text{Cov}(X_{n},F_{2})$
From the true model:
$X_{n}=\gamma_{n}Z+\beta_{n}F_{2}+\varepsilon_{n}\quad\Rightarrow\quad\text{Cov}(X_{n},F_{2})=\gamma_{n}\cdot\text{Cov}(Z,F_{2})+\beta_{n}\cdot\text{Var}(F_{2})=\gamma_{n}\delta+\beta_{n}$
Thus:
$\hat{\beta}_{n}=\beta_{n}+\gamma_{n}\delta$
This shows that omitting the confounder $Z$ introduces systematic bias into the estimated factor loading $\hat{\beta}_{n}$, depending on the true confounder exposure $\gamma_{n}$ and its correlation $\delta$ with $F_{2}$ (López de Prado et al.,, 2024).
3.1.2 Resulting Portfolio Inefficiency
Suppose an investor constructs a minimum-variance portfolio targeting exposure $c^{\top}=[0,1]$ using the misspecified loadings:
$B_{\text{miss}}=\begin{bmatrix}\hat{\beta}_{1}\\ \hat{\beta}_{2}\end{bmatrix}=\begin{bmatrix}\beta_{1}+\gamma_{1}\delta\\ \beta_{2}+\gamma_{2}\delta\end{bmatrix}$
The optimizer solves:
$\omega^{*}_{\text{miss}}=B_{\text{miss}}^{\top}{}^{-1}c^{\top}=\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\begin{bmatrix}-1\\ 1\end{bmatrix}$
We evaluate the true exposure delivered by this portfolio:
$\omega^{*}_{\text{miss}}{}^{\top}B_{\text{true}}=\left(\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\begin{bmatrix}-1&1\end{bmatrix}\right)\begin{bmatrix}\gamma_{1}&\beta_{1}\\ \gamma_{2}&\beta_{2}\end{bmatrix}=\frac{1}{\hat{\beta}_{2}-\hat{\beta}_{1}}\left(\gamma_{2}-\gamma_{1},\ \beta_{2}-\beta_{1}\right)$
This actual exposure deviates from the intended $[0,1]$ unless $\gamma_{1}=\gamma_{2}=0$, i.e., no confounding. The portfolio is therefore inefficient even though the investor used valid optimization logic on misspecified inputs. In summary, from (López de Prado et al.,, 2024, 2025) even when expected returns and covariances are correctly specified, factor misspecification alone is sufficient to induce portfolio inefficiency. This does not require causal reasoning, it arises purely from statistical misspecification, counter to arguments made in (López de Prado et al.,, 2025).
3.1.3 Conditions Underpinning the Necessity Claim
The central thesis of (López de Prado et al.,, 2025) is that causal factor modeling is a necessary condition for investment efficiency. However, this claim rests on several assumptions, some explicit, others implicit, which are listed and critically evaluated in Table 1.
Condition Description True/False
C1 Causal factor models are necessary for portfolio efficiency. False — disproven by counterexample.
C2 Associational models (e.g., correlation-based) are always misspecified due to confounders or colliders. False — not true if omitted variables are independent.
C3 Mean-variance optimization requires unbiased exposures $B$. False — other allocation rules exist (e.g., classification-based).
C4 The data-generating process is linear and additive. False — unrealistic in many real-world markets.
C5 All observed associations must be interpreted causally; correlation cannot justify predictive modeling. False — contradicted by predictive theory (see (Shmueli,, 2010)).
C6 Causal discovery tools can recover valid DAGs from financial data. Uncertain — relies on strong assumptions (e.g., no latent confounders, i.i.d.).
C7 The true causal graph is stable and estimable. False — often untrue due to nonstationarity and market adaptation.
Table 1: Conditions required for the necessity of causal modeling, as claimed in (López de Prado et al.,, 2025), and prior work (López de Prado et al.,, 2024; López de Prado and Zoonekynd,, 2024).
As shown in Table 1, the necessity claim relies on several assumptions that often do not hold in practical or general settings. If even one of these conditions fails, causal factor modeling cannot be asserted as a general requirement for investment efficiency. As our counterexamples and derivations show, Condition C1 is refuted outright, while C2–C5 are context-dependent and frequently violated in financial modeling.
The following Section introduces the structural cancellation counterexample, in which omitted variables are neutralized through bias absorption.
3.2 Structural Cancellation as a Counterexample
We begin with a counterexample showing that omitted variables do not inevitably cause signal inversion. Instead, omitted-variable bias can produce approximate cancellation, preserving directional validity while attenuating magnitudes. This challenges prior claims that structural misspecification necessarily yields inverted signals and collapsed frontiers.
3.2.1 Setup
Let the observed feature be $X\sim\mathcal{N}(0,1)$ and the latent confounder be
$Z^{\prime}=-\alpha X+\eta+\zeta,\quad\eta\sim\mathcal{N}(0,\sigma_{\eta}^{2}),\quad\zeta\sim\mathcal{N}(0,\sigma_{\zeta}^{2}),$
independent of $X$. The true signal is
$s_{\mathrm{true}}=\beta X+\gamma Z^{\prime},$
and returns are generated as
$Y=s_{\mathrm{true}}+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,\sigma^{2}).$
3.2.2 Misspecified Estimation
Suppose we regress $Y$ on $X$ alone, omitting $Z^{\prime}$. The OLS coefficient is
$\hat{\beta}=\frac{\mathbb{E}[XY]}{\mathbb{E}[X^{2}]}.$
Substituting,
$\hat{\beta}=\frac{\mathbb{E}[X(\beta X+\gamma Z^{\prime}+\varepsilon)]}{\mathbb{E}[X^{2}]}=\beta+\gamma\cdot\frac{\mathbb{E}[XZ^{\prime}]}{\mathbb{E}[X^{2}]}.$
Because $Z^{\prime}=-\alpha X+\eta+\zeta$, we obtain
$\mathbb{E}[XZ^{\prime}]=-\alpha\,\mathbb{E}[X^{2}]=-\alpha.$
Thus,
$\hat{\beta}=\beta-\alpha\gamma.$
Lemma 1 (Structural Cancellation).
If $\beta$ and $\alpha\gamma$ have the same sign and comparable magnitude, the OLS coefficient
$\hat{\beta}=\beta-\alpha\gamma$
remains positive but attenuated relative to $\beta$. Hence, omitted-variable bias may shrink weights toward zero rather than invert them.
Proof.
If $\beta>0$ and $\alpha\gamma>0$, then $\hat{\beta}=\beta-\alpha\gamma$. When $0<\alpha\gamma<\beta$, we have $0<\hat{\beta}<\beta$. Thus, directionality is preserved, though magnitude is reduced. An analogous argument applies for $\beta<0$. ∎
Corollary 1 (Omitted-Variable Bias Does Not Necessarily Imply Inefficiency).
Under the conditions of Lemma 1, the misspecified estimator $\hat{\beta}$ retains the correct sign. Therefore, directional informativeness is preserved, and the resulting portfolio weights remain viable for convex frontier construction. Misspecification reduces efficiency through attenuation but does not induce systematic failure.
Remark. Contrary to prior claims, sign inversion requires $\alpha\gamma>\beta$, not misspecification per se. In practice, moderate confounding is more likely to yield attenuated but still directionally valid signals.
Figure 1 shows a Monte Carlo simulation of the cancellation effect. Despite omitting $Z^{\prime}$, the estimated coefficients align in sign with the true signal but exhibit attenuation. This demonstrates that misspecification can preserve frontier validity when structural dependence partially cancels.
3.3 Optimization Under Misspecification
We now distinguish prediction, ranking, and efficient allocation. Our results demonstrate that directional agreement between signals is necessary but not sufficient for portfolio efficiency: signals may align in sign yet still yield inefficient allocations due to miscalibration. This refines prior claims that inefficiency arises only through sign inversion.
Theorem 1 (Directional Agreement Does Not Imply Efficiency).
There exist predictive signals $\omega_{\mathrm{pred}}$ such that
$\operatorname{sign}(\omega_{\mathrm{pred}}(x))=\operatorname{sign}(\omega_{\mathrm{true}}(x))\quad\forall x\in\mathcal{X},$
yet
$R(\omega_{\mathrm{pred}})\neq R(\omega_{\mathrm{true}}),\quad\mathrm{Sharpe}(\omega_{\mathrm{pred}})<\mathrm{Sharpe}(\omega_{\mathrm{true}}),$
where $R(\omega)=\mathbb{E}[\omega(x)r(x)]$ and $\mathrm{Sharpe}(\omega)=\tfrac{\mathbb{E}[\omega(x)r(x)]}{\sqrt{\mathrm{Var}(\omega(x)r(x))}}$.
Proof.
Let $\omega_{\mathrm{pred}}(x)=\epsilon\,\omega_{\mathrm{true}}(x)$ with $\epsilon\in(0,1)$. Then
$R(\omega_{\mathrm{pred}})=\epsilon R(\omega_{\mathrm{true}}),\quad\mathrm{Sharpe}(\omega_{\mathrm{pred}})=\epsilon\,\mathrm{Sharpe}(\omega_{\mathrm{true}}).$
Directionality is preserved, but performance is strictly attenuated when $\epsilon\neq 1$. The case $\epsilon<0$ recovers full inversion, a special case of inefficiency previously emphasized. ∎
Remark. Directional correctness ensures that signs match, but without proper scaling, efficiency contracts.
Lemma 2 (Ranking Does Not Guarantee Efficient Sizing).
Let $f_{\mathrm{true}},f_{\mathrm{pred}}:\mathcal{X}\to\mathbb{R}$ define weights $\omega_{\mathrm{true}}(x)=f_{\mathrm{true}}(x)$, $\omega_{\mathrm{pred}}(x)=f_{\mathrm{pred}}(x)$. Suppose rankings agree:
$f_{\mathrm{true}}(x_{i})>f_{\mathrm{true}}(x_{j})\iff f_{\mathrm{pred}}(x_{i})>f_{\mathrm{pred}}(x_{j}).$
If $f_{\mathrm{pred}}$ is not a positive multiple of $f_{\mathrm{true}}$, then
$\mathbb{E}\big{[}(\omega_{\mathrm{pred}}(x)-\omega_{\mathrm{true}}(x))^{2}\big{]}>0,$
so efficiency is lost through miscalibration.
Proof.
If $f_{\mathrm{pred}}=af_{\mathrm{true}}$ for some $a>0$, weights differ only by scale and efficiency is preserved. Otherwise, the deviation $\delta(x)=f_{\mathrm{pred}}(x)-f_{\mathrm{true}}(x)$ is nonzero with positive probability, hence $\mathbb{E}[\delta(x)^{2}]>0$. This implies inefficient sizing under mean–variance optimization. ∎
Corollary 2 (Calibration Is Necessary for Efficiency).
Even if signals preserve direction and ranking, portfolios formed from miscalibrated magnitudes are strictly less efficient than those based on correctly scaled signals.
Proof.
Follows directly from Lemma 2: unless $f_{\mathrm{pred}}=af_{\mathrm{true}}$ with $a>0$, deviations in magnitude reduce realized efficiency. ∎
Remark. Ranking ensures membership in the efficient set, but calibration determines the investor’s position along the frontier.
Proposition 1 (Sharpe Ratio with Surrogate Signals).
Let $\mu$ be the true mean return vector, $\tilde{\mu}$ a surrogate preserving its order, and $V\succ 0$ the covariance matrix. Then
$\mathrm{Sharpe}(V^{-1}\tilde{\mu})=\frac{\mu^{\top}V^{-1}\tilde{\mu}}{\sqrt{\tilde{\mu}^{\top}V^{-1}\tilde{\mu}}}=\|\mu\|_{V^{-1}}\cdot\rho_{V^{-1}},$
where $\rho_{V^{-1}}=\tfrac{\mu^{\top}V^{-1}\tilde{\mu}}{\|\mu\|_{V^{-1}}\|\tilde{\mu}\|_{V^{-1}}}\in[-1,1]$.
Proof.
By Almgren and Chriss, (2005), order preservation ensures $V^{-1}\tilde{\mu}$ lies in the efficient cone. Computing return and variance under $V^{-1}$ geometry yields the stated scaling law. ∎
Lemma 3 (Ranking–Calibration Decomposition).
Ranking is necessary for efficient-set membership, while calibration governs quantitative efficiency within the set. Formally: if ranking fails, efficiency is unattainable; if ranking holds, Sharpe scales linearly with $\rho_{V^{-1}}$.
Proof.
Necessity: if ordering is violated, $\tilde{\mu}$ may exit the efficient cone, making optimization invalid. Sufficiency: if ranking holds, Proposition 1 shows $\mathrm{Sharpe}(V^{-1}\tilde{\mu})=\|\mu\|_{V^{-1}}\rho_{V^{-1}}$, so calibration (alignment $\rho_{V^{-1}}$) determines efficiency. ∎
Remark. Directional correctness is necessary but insufficient; ranking secures frontier validity, and calibration dictates realized Sharpe. This decomposition explains why misspecified models underperform via attenuation rather than systematic inversion.
3.4 Testing Optimization Geometry Under Misspecification
We now examine whether structural misspecification necessarily collapses portfolio optimization geometry. Using nonlinear confounded processes, we show that surrogate signals, even when estimated from misspecified models, can yield viable efficient frontiers provided they remain directionally aligned with the true return vector.
3.4.1 Nonlinear Confounded Signals
Consider observable features $X_{1},X_{2}\sim\mathcal{N}(0,1)$ and an unobserved confounder
$Z=\alpha_{1}X_{1}+\alpha_{2}X_{2}+\eta,\quad\eta\sim\mathcal{N}(0,1).$
Let the true signal follow the nonlinear model
$\mathbb{P}(Y=1\mid X,Z)=\sigma\!\big{(}\tanh(X)+0.5\sin(Z)\big{)},$
with portfolio weights
$\omega^{\mathrm{true}}=2\,\mathbb{P}(Y=1\mid X,Z)-1.$
A misspecified logistic regression omitting $Z$ produces surrogate signals
$\omega^{\mathrm{pred}}=2\,\sigma(\hat{\beta}^{\top}X)-1.$
This setup extends the structural cancellation mechanism: omitted-variable bias and nonlinearity need not eliminate directional informativeness.
3.4.2 Robustness of Optimization Geometry
We formalize the claim that directional alignment, not structural fidelity, secures valid frontiers.
Lemma 4 (Directional Signals Yield Valid Frontiers).
Let $\mu$ denote true expected returns, $\tilde{\mu}$ surrogate expected returns, and $\Sigma\succ 0$ the covariance matrix. Consider the Markowitz problem
$\max_{\omega\in\mathbb{R}^{n}}\quad\tilde{\mu}^{\top}\omega-\lambda\,\omega^{\top}\Sigma\omega\quad\text{s.t. }\mathbf{1}^{\top}\omega=1,\ \lambda>0.$
If $\tilde{\mu}^{\top}\mu>0$, then the efficient frontier traced by
$\mathcal{F}(\lambda)=\big{(}\sigma(\omega^{*}(\lambda)),\,\mu^{\top}\omega^{*}(\lambda)\big{)}$
is smooth, convex, and non-degenerate in the $(\sigma,\mathbb{E}[r])$ plane.
Proof.
The Lagrangian first-order condition gives
$\omega^{*}(\lambda)=\tfrac{1}{2\lambda}\Sigma^{-1}(\tilde{\mu}-\nu\mathbf{1}),$
with $\nu$ uniquely determined by the budget constraint. Since $\Sigma\succ 0$, $\omega^{*}(\lambda)$ is unique and varies smoothly with $\lambda$. The realized return is
$R(\lambda)=\mu^{\top}\omega^{*}(\lambda)\propto\mu^{\top}\Sigma^{-1}\tilde{\mu}$.
If $\tilde{\mu}^{\top}\mu>0$, then $R(\lambda)>0$ for all $\lambda>0$. Portfolio variance $\sigma^{2}(\lambda)=\omega^{*\top}\Sigma\omega^{*}$ is strictly convex in $\lambda$, so the mapping $\lambda\mapsto(\sigma(\lambda),R(\lambda))$ is smooth and convex. Non-degeneracy follows from $R(\lambda)>0$. ∎
Corollary 3 (Frontier Preservation Under Misspecification).
Suppose $\tilde{\mu}$ is estimated from a misspecified logistic regression omitting confounders. If $\tilde{\mu}^{\top}\mu>0$, then the Markowitz frontier remains smooth, convex, and non-degenerate.
Proof.
Immediate from Lemma 4, with $\tilde{\mu}$ interpreted as a surrogate mean vector. Directional alignment ensures positive realized returns, convexity, and smooth variation of $\omega^{*}(\lambda)$. ∎
Remark. Structural misspecification does not collapse optimization geometry. Directional informativeness suffices to preserve frontier viability.
3.4.3 Sharpe Ratio Sensitivity to Alignment
We next quantify how efficiency contracts as alignment weakens.
Proposition 2 (Sharpe Ratio Scales with Alignment).
Let $\mu\in\mathbb{R}^{n}$ denote true returns and $\nu\perp\mu$ with $\|\nu\|=\|\mu\|$. Define
$\tilde{\mu}(\theta)=\cos\theta\,\mu+\sin\theta\,\nu,\quad\theta\in[0,\pi].$
For $\omega^{*}(\theta)=\tfrac{1}{2\lambda}\Sigma^{-1}\tilde{\mu}(\theta)$, the realized Sharpe ratio satisfies
$\mathrm{Sharpe}(\omega^{*}(\theta))=\cos\theta\cdot\mathrm{Sharpe}(\omega^{*}(0)).$
Proof.
We have
$\mu^{\top}\omega^{*}(\theta)=\tfrac{1}{2\lambda}\mu^{\top}\Sigma^{-1}\tilde{\mu}(\theta)=\cos\theta\cdot\mu^{\top}\omega^{*}(0),$
since $\mu^{\top}\Sigma^{-1}\nu=0$. Portfolio volatility $\sigma(\omega^{*}(\theta))=\tfrac{1}{2\lambda}\sqrt{\tilde{\mu}(\theta)^{\top}\Sigma^{-1}\tilde{\mu}(\theta)}$ is invariant in $\theta$ by construction. Dividing return by volatility yields the stated scaling. ∎
Corollary 4 (Frontier Collapse Under Misalignment).
If $\cos\theta\leq 0$, then $\mathrm{Sharpe}(\omega^{*}(\theta))\leq 0$. Thus:
  • $\cos\theta=0$: the portfolio is orthogonal to $\mu$ and achieves zero return.
  • $\cos\theta<0$: the portfolio is anti-aligned and delivers negative performance.
Positive directional alignment is therefore necessary for viable optimization.
Remark. Efficiency degrades linearly with cosine alignment $\rho=\cos\theta$. Collapse occurs only at $\rho\leq 0$, confirming that miscalibration, not inversion, is the dominant inefficiency channel.
3.5 Approximate Cancellation and Attenuation
We study omitted‐variable bias when a confounder nearly cancels the observed signal but is corrupted by independent noise. Let $X\sim\mathcal{N}(0,1)$, and let
$Z^{\prime}\;=\;-\alpha\,X+\eta+\zeta,\qquad\eta\sim\mathcal{N}(0,\sigma_{\eta}^{2}),\ \ \zeta\sim\mathcal{N}(0,\sigma_{\zeta}^{2}),$
with $\eta,\zeta$ independent of $X$ and of each other. Define the normalized confounder channel
$\widetilde{Z}\;:=\;\frac{Z^{\prime}}{\sqrt{1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2}}},$
so that $\operatorname{Var}(\widetilde{Z})=\operatorname{Var}(Z^{\prime})/(1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2})=1$. Consider the structural outcome
$Y\;=\;\beta\,X+\gamma\,\widetilde{Z}+\varepsilon,\qquad\mathbb{E}[\varepsilon\mid X,\widetilde{Z}]=0,\ \ \operatorname{Var}(\varepsilon)=\sigma_{\varepsilon}^{2},$
and suppose a practitioner fits a misspecified regression of $Y$ on $X$ only. Denote by $\widehat{\beta}_{\text{OLS}}(\sigma_{\zeta})$ the population OLS coefficient of $Y$ on $X$ obtained under this misspecification when the confounding‐noise level is $\sigma_{\zeta}$.
The normalization of $Z^{\prime}$ into $\widetilde{Z}$ is natural in practice: the structural channel (the loading $\gamma$) acts on a confounder with unit variance. As the extraneous noise $\zeta$ increases, the correlation between $X$ and $\widetilde{Z}$ attenuates smoothly, producing a continuous bias‐decay law that we derive exactly below.
Proposition 3 (Exact attenuation law under approximate cancellation).
With the setup above (and $\operatorname{Var}(X)=1$), the misspecified OLS slope is
$\widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})\;=\;\beta\;-\;\frac{\alpha\,\gamma}{\sqrt{1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2}}}.$ (1)
In particular, as the confounding noise increases,
$\lim_{\sigma_{\zeta}\downarrow 0}\widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})\;=\;\beta\;-\;\frac{\alpha\gamma}{\sqrt{1+\sigma_{\eta}^{2}}},\qquad\lim_{\sigma_{\zeta}\uparrow\infty}\widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})\;=\;\beta,$
and $\widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})$ is strictly increasing in $\sigma_{\zeta}$ whenever $\alpha\gamma>0$.
Proof.
Since $\operatorname{Var}(X)=1$, population OLS gives $\widehat{\beta}_{\text{OLS}}=\frac{\operatorname{Cov}(X,Y)}{\operatorname{Var}(X)}=\operatorname{Cov}(X,Y).$ Using $Y=\beta X+\gamma\widetilde{Z}+\varepsilon$ and $\mathbb{E}[\varepsilon\mid X,\widetilde{Z}]=0$,
$\operatorname{Cov}(X,Y)=\beta\,\operatorname{Var}(X)+\gamma\,\operatorname{Cov}(X,\widetilde{Z})=\beta+\gamma\,\operatorname{Cov}\!\Big{(}X,\frac{Z^{\prime}}{\sqrt{1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2}}}\Big{)}.$
Because $Z^{\prime}=-\alpha X+\eta+\zeta$ with $\eta,\zeta$ independent of $X$, $\operatorname{Cov}(X,Z^{\prime})=-\alpha\,\operatorname{Var}(X)=-\alpha.$ Hence
$\operatorname{Cov}(X,\widetilde{Z})=\frac{\operatorname{Cov}(X,Z^{\prime})}{\sqrt{1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2}}}=-\frac{\alpha}{\sqrt{1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2}}}.$
Substituting yields the formula in the statement. The limits as $\sigma_{\zeta}$ tends to 0 or $\infty$ are immediate, and monotonicity for $\alpha\gamma>0$ follows from differentiating w.r.t. $\sigma_{\zeta}$:
$\frac{\partial\widehat{\beta}_{\text{OLS}}}{\partial\sigma_{\zeta}}=\alpha\gamma\,\frac{\sigma_{\zeta}}{\big{(}1+\sigma_{\eta}^{2}+\sigma_{\zeta}^{2}\big{)}^{3/2}}\ \geq\ 0,$
with strict positivity for $\sigma_{\zeta}>0$. ∎
Corollary 5 (Continuity, smoothness, and bounds).
The map $\sigma_{\zeta}\mapsto\widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})$ is real‐analytic on $0,\infty)$. Moreover,
$\beta-\frac{|\alpha\gamma|}{\sqrt{1+\sigma_{\eta}^{2}}}\ \leq\ \widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})\ \leq\ \beta,\qquad\forall\,\sigma_{\zeta}\geq 0,$
with the lower (upper) bound attained at $\sigma_{\zeta}=0$ ($\sigma_{\zeta}\to\infty$) when $\alpha\gamma>0$.
Proof.
Real‐analyticity follows because $\widehat{\beta}_{\text{OLS}}$ is a composition of analytic functions of $\sigma_{\zeta}$. The bounds are the minimum/maximum of the strictly increasing function in Proposition 3 (for $\alpha\gamma>0$); the reversed inequalities hold when $\alpha\gamma<0$. ∎
Corollary 6 (Non‐inversion under approximate cancellation).
If $\beta>0$ and $\alpha\gamma\leq\beta\sqrt{1+\sigma_{\eta}^{2}}$, then
$\widehat{\beta}_{\emph{OLS}}(\sigma_{\zeta})\ \geq\ \beta-\frac{\alpha\gamma}{\sqrt{1+\sigma_{\eta}^{2}}}\ \geq\ 0\qquad\forall\,\sigma_{\zeta}\geq 0.$
Hence sign inversion cannot occur for any $\sigma_{\zeta}\geq 0$. Analogous conditions hold for $\beta<0$.
Proof.
Combine the lower bound in Corollary 5 with the assumption $\alpha\gamma\leq\beta\sqrt{1+\sigma_{\eta}^{2}}$ when $\beta>0$. ∎
Implications for signal alignment and Sharpe.
Let $w(\nu)\propto V^{-1}\nu$ denote the mean–variance direction for a signal $\nu$, and evaluate performance under the true mean $\mu$. For any surrogate $\tilde{\mu}$,
[…]
with $\rho_{V^{-1}}\in[-1,1]$. If $\tilde{\mu}(\sigma_{\zeta})=a(\sigma_{\zeta})\,\mu+b(\sigma_{\zeta})\,\nu$ where $\nu\perp_{V^{-1}}\mu$ and $a,b$ are continuous in $\sigma_{\zeta}$ (as induced by the scalar attenuation in Proposition 3), then $\rho_{V^{-1}}(\mu,\tilde{\mu}(\sigma_{\zeta}))=\frac{a(\sigma_{\zeta})}{\sqrt{a(\sigma_{\zeta})^{2}+b(\sigma_{\zeta})^{2}}}$ is continuous and strictly increasing in any regime where $a(\sigma_{\zeta})$ increases while $b(\sigma_{\zeta})$ is weakly decreasing. Consequently, $\mathrm{Sharpe}(V^{-1}\tilde{\mu}(\sigma_{\zeta}))$ varies continuously with $\sigma_{\zeta}$ and contracts smoothly toward zero alignment as $a(\sigma_{\zeta})\downarrow 0$ (or expands as $a(\sigma_{\zeta})\uparrow 1$).
Proposition 4 (Continuity and Lipschitz control of Sharpe under attenuation).
Suppose $\tilde{\mu}(\sigma_{\zeta})$ is constructed by attenuating the confounding component as in Proposition 3 (componentwise, with independent channels) and $\|\tilde{\mu}(\sigma_{\zeta})\|_{V^{-1}}$ is uniformly bounded away from zero. Then
$\Big{|}\mathrm{Sharpe}\big{(}V^{-1}\tilde{\mu}(\sigma_{\zeta})\big{)}-\mathrm{Sharpe}\big{(}V^{-1}\tilde{\mu}(\sigma_{\zeta}^{\prime})\big{)}\Big{|}\ \leq\ L\,|\sigma_{\zeta}-\sigma_{\zeta}^{\prime}|$
for some $L<\infty$ depending on $(\mu,V^{-1})$ and the attenuation profiles; in particular, the Sharpe function is continuous (indeed locally Lipschitz) in $\sigma_{\zeta}$.
Proof.
Write $S(\sigma)=\frac{\mu^{\top}V^{-1}\tilde{\mu}(\sigma)}{\sqrt{\tilde{\mu}(\sigma)^{\top}V^{-1}\tilde{\mu}(\sigma)}}$. By the quotient rule and Cauchy–Schwarz in the $V^{-1}$ inner product,
$|S^{\prime}(\sigma)|\ \leq\ \frac{\|\mu\|_{V^{-1}}\ \|\tilde{\mu}^{\prime}(\sigma)\|_{V^{-1}}}{\|\tilde{\mu}(\sigma)\|_{V^{-1}}}\;+\;\frac{\|\mu\|_{V^{-1}}\ \|\tilde{\mu}(\sigma)\|_{V^{-1}}\ \|\tilde{\mu}^{\prime}(\sigma)\|_{V^{-1}}}{\|\tilde{\mu}(\sigma)\|_{V^{-1}}^{3}}\ \leq\ C\,\|\tilde{\mu}^{\prime}(\sigma)\|_{V^{-1}},$
for a finite constant $C$ whenever $\|\tilde{\mu}(\sigma)\|_{V^{-1}}$ is bounded away from zero. The attenuation map $\sigma\mapsto\tilde{\mu}(\sigma)$ is smooth in each component (by Proposition 3), hence $\|\tilde{\mu}^{\prime}(\sigma)\|_{V^{-1}}$ is bounded on compact intervals, giving the Lipschitz bound by the mean value theorem. ∎
[…]
4 Numerical Results
We now provide empirical validation of the theoretical framework developed in Section 3, focusing on the geometry of portfolio optimization under structural misspecification. The experiments are organized in five parts. First, simulations illustrate structural cancellation, showing that omitted-variable bias may attenuate but not invert signals (Corollary 1). Second, calibration–ranking experiments validate that ranking ensures efficient-set membership, whereas calibration governs quantitative efficiency (Lemma 3, Proposition 1). Third, nonlinear counterexamples test robustness of optimization geometry, demonstrating that misspecification does not collapse the frontier (Lemma 4, Corollary 3). Fourth, alignment experiments confirm that Sharpe ratios contract linearly with cosine alignment, as predicted by Proposition 2 and Corollary 4. Finally, empirical validation with S&P 500 equity data shows that predictive signals from misspecified models can sustain convex and coherent frontiers in practice.
Together, these results disentangle the distinct roles of directionality, ranking, and calibration in determining portfolio outcomes. All simulation code is available in the supplementary material.
4.1 Structural Cancellation Under Misspecification
We simulate $n=1000$ observations with features $X\sim\mathcal{N}(0,1)$ and define a confounder $Z^{\prime}=-\alpha X+\eta$, where $\alpha=0.6$ and $\eta\sim\mathcal{N}(0,1-\alpha^{2})$. Outcomes depend on both $X$ and $Z^{\prime}$, but the predictive model uses only $X$. Portfolio weights are computed as
$\omega^{\mathrm{true}}=2\cdot\mathbb{P}(Y=1\mid X,Z^{\prime})-1,\quad\omega^{\mathrm{pred}}=2\cdot\mathbb{P}(Y=1\mid X)-1.$
Figure 1 shows strong directional alignment without sign inversions, confirming Lemma 1. Magnitudes are attenuated, but the frontier remains viable, illustrating Corollary 1.
Figure 1: Portfolio Weights under Structural Cancellation. Scatter of true weights (x-axis) against misspecified weights (y-axis). Signs are preserved but magnitudes are attenuated, showing directional consistency despite misspecification.
4.2 Calibration and Ranking within the Efficient Set
To isolate calibration effects, we generate synthetic return vectors $\mu\in\mathbb{R}^{n}$ and construct surrogate signals $\tilde{\mu}$ via monotone transformations:
$\tilde{\mu}_{i}=\operatorname{sign}(\mu_{i})\,|\mu_{i}|^{p},\quad p\in[0.6,1.4].$
These transformations preserve ranking (Lemma 3), but distort magnitudes. Panel (a) of Figure 2 shows order preservation; Panel (b) demonstrates that Sharpe efficiency peaks at $p=1$ and contracts smoothly otherwise, confirming Proposition 1. Inefficiency thus stems primarily from calibration errors, not directional failure.
(a) Ranking preserved under monotone transforms ($p=0.6,0.8,1.0$). Order is maintained across assets, ensuring efficient-set membership.
(b) Sharpe ratios as a function of $p$. Efficiency peaks at $p=1$ and declines smoothly for $p\neq 1$, consistent with Proposition 1.
Figure 2: Calibration and Ranking Experiments. Monte Carlo validation of Section $3.$ Panel (a): monotone transforms preserve ranking, consistent with Lemma 3. Panel (b): Sharpe efficiency contracts with miscalibration, confirming Proposition 1. Together, these results show that ranking ensures validity, while calibration determines frontier position.
4.3 Nonlinear Counterexamples and Frontier Geometry
We next consider the nonlinear confounded model of Section 3.4. True and misspecified signals remain directionally aligned (Figure 3), contradicting the claim that omitted variables necessarily induce inversion. Moving to full portfolios, Figure 4 shows that efficient frontiers constructed from misspecified signals remain smooth and convex, empirically validating Lemma 4 and Corollary 3.
Figure 3: Misspecified vs. True Portfolio Weights. Scatter of predicted versus true weights under a nonlinear confounded model. Misspecification compresses magnitudes but does not induce systematic sign inversion.
Figure 4: Efficient Frontier from Misspecified Signals. Portfolios based on misspecified signals yield a smooth, convex frontier, illustrating preserved geometry despite reduced efficiency.
4.4 Sensitivity to Alignment
Using surrogate vectors $\tilde{\mu}(\theta)=\cos(\theta)\mu+\sin(\theta)\nu$, we test efficiency as alignment $\rho=\cos(\theta)$ varies. Figure 5 shows (a) frontiers contract proportionally to $\rho$ and (b) Sharpe ratios scale linearly, exactly as predicted by Proposition 2 and Corollary 4. These experiments demonstrate that alignment is the critical determinant of efficiency.
(a) Frontiers contract as $\rho$ decreases.
(b) Sharpe ratio scales linearly with $\rho$.
Figure 5: Sensitivity to Alignment. (a) Efficient frontiers shrink smoothly as alignment $\rho$ declines. (b) Sharpe ratios scale linearly with $\rho$, confirming the theoretical prediction.
4.5 Empirical Validation on Equity Data
Finally, we apply the framework to daily returns of five S&P 500 stocks over 1000 trading days. Logistic predictors trained on lagged features and volatility measures provide signals $\omega_{j}$, which feed into a quadratic optimizer. Figure 6 shows that the resulting frontier is smooth and convex despite structural misspecification, confirming that directional validity and approximate calibration suffice for practical viability.
(a) Mean–volatility scatter.
(b) Empirical frontier.
Figure 6: Real-World Signals from S&P 500 Stocks. (a) Scatter of assets by mean and volatility. (b) Empirical efficient frontier remains smooth and convex, showing that directional signals yield coherent portfolios even under misspecification.
4.6 Monte Carlo Validation of Bias Attenuation
Varying the confounding noise parameter $\sigma_{\zeta}$, we regress $Y=\beta X+\gamma\widetilde{Z}+\varepsilon$ on $X$ alone. Figure 7 shows empirical OLS estimates align with the theoretical attenuation curve from Equation (1). This confirms Corollary 1: omitted-variable bias degrades signals smoothly, not catastrophically.
(a) Monte Carlo estimates vs. theory.
$\sigma_{\zeta}$ $\hat{\beta}_{\mathrm{OLS}}$ $\beta_{\mathrm{theory}}$
0.000 -0.2203 -0.2200
0.050 -0.2187 -0.2187
0.100 -0.2163 -0.2163
0.200 -0.2099 -0.2099
0.300 -0.2020 -0.2020
0.400 -0.1929 -0.1929
0.500 -0.1828 -0.1829
0.600 -0.1720 -0.1720
0.700 -0.1605 -0.1606
0.800 -0.1486 -0.1486
(b) Selected numerical values.
Figure 7: Attenuation of Omitted-Variable Bias. (a) Monte Carlo simulations (dots) align with the theoretical curve (dashed). (b) Numerical estimates match analytical predictions across noise levels, showing that omitted-variable bias diminishes smoothly as $\sigma_{\zeta}$ increases.
5 Discussion
This paper has revisited the claim that causal factor models are a prerequisite for portfolio optimization. Through theoretical results, simulation counterexamples, and empirical validation, we show that this assertion does not hold in general. Even when structurally misspecified, predictive models can produce signals that remain directionally aligned with true returns and sufficient to sustain smooth, convex efficient frontiers.
Our analysis clarifies several distinctions often blurred in recent debates. First, we separate directionality, ranking, and calibration. Theorem 1, Lemma 2, and Corollary 2 establish that directional correctness and ranking are necessary but not sufficient for efficiency: miscalibration attenuates realized returns and Sharpe ratios. This explains why predictive accuracy in classification does not automatically translate into optimization efficiency, echoing the explanatory–predictive divide highlighted by Shmueli, (2010).
Second, we refine the interplay between ranking and calibration. In the sort-based framework of Almgren and Chriss, (2005), correct ranking ensures Pareto efficiency. Proposition 1 and Lemma 3 extend this by showing that calibration governs quantitative performance within the efficient set. Figures 2 and 5 confirm empirically that while ranking secures validity, calibration determines realized efficiency.
Third, we counter the claim that misspecification necessarily collapses optimization geometry. Prior critiques (e.g. (López de Prado et al.,, 2024, 2025)) argued that omitted variables or nonlinearities produce inverted signals and degenerate frontiers. Our counterexamples in Sections 3.2 and 3.4 show the opposite: predictive signals often retain directional informativeness, and frontiers remain convex and non-degenerate.
Fourth, Section 3.5 demonstrates that signal viability is not knife-edge. Both analytically and empirically, omitted-variable bias attenuates continuously as dependence weakens (Figure 7). This continuity result highlights robustness: efficiency degrades smoothly rather than collapsing.
Finally, our empirical study on S&P 500 equities (Figure 6) shows that predictive signals yield coherent frontiers in practice. Even without structural correctness, directional validity with approximate calibration suffices for actionable portfolio design. Predictive modeling should therefore be viewed as a complementary and often more pragmatic tool when structural assumptions are uncertain, unstable, or untestable.
While causal inference remains valuable for attribution, regime analysis, and risk decomposition (Lo,, 2017; Wilcox and Gebbie,, 2014; Rodriguez Dominguez,, 2023), our results demonstrate that causal identifiability is not a prerequisite for optimization efficiency. The decisive benchmarks are whether signals preserve directional informativeness, maintain correct ranking, and achieve sufficient calibration. By formalizing this decomposition and validating it empirically, we provide a constructive correction to structural critiques and advocate for evaluation criteria that recognize the robustness and operational utility of predictive-first approaches.
6 Conclusion and Future Work
We have offered a systematic re-evaluation of the claim that causal factor models are necessary for portfolio efficiency. Through formal analysis, controlled simulations, and empirical evidence, we show that structural misspecification, such as omitted variables or nonlinear confounding, does not inherently invalidate optimization. Predictive models without causal identifiability can nevertheless yield directionally valid signals that sustain smooth and convex efficient frontiers.
Our contributions are fourfold. First, we disentangle directionality, ranking, and calibration, proving that directional correctness is necessary but insufficient when magnitudes are miscalibrated. Second, we show that structurally misspecified yet directionally aligned signals preserve optimization geometry, extending the portfolios-from-sorts framework by Almgren and Chriss, (2005) to account for calibration effects. Third, we demonstrate analytically and empirically that signal viability is robust, with omitted-variable bias attenuating smoothly rather than collapsing abruptly. Fourth, we validate these findings with real market data, confirming that predictive signals yield coherent frontiers even absent strict causal grounding.
These results caution against overstating the necessity of causal inference in portfolio design. While causal models remain indispensable for attribution, risk decomposition, and policy analysis (Lo,, 2017; Wilcox and Gebbie,, 2014; Rodriguez Dominguez,, 2023), their superiority is not automatic. What ultimately governs optimization performance is the directional fidelity and calibration of predictive signals, not structural correctness per se. This perspective motivates hybrid approaches that integrate causal diagnostics with predictive-first models to balance interpretability and robustness.
Future work should pursue three directions: integrating causal diagnostics directly into predictive pipelines, systematically benchmarking causal versus predictive models under regime shifts, and extending misspecification-resilient optimization theory to high-dimensional and nonlinear contexts. Such research can clarify when a causal structure enhances decision-making and when predictive validity alone is sufficient.
To sum up, structural correctness may be desirable, but it is not a prerequisite for portfolio efficiency. Effective portfolio construction relies on signals that are directionally valid, properly calibrated, and resilient to the inevitable imperfections inherent in financial modeling.
Appendix A Simulation Protocols and Code
All simulations were implemented in Python using NumPy, Matplotlib, and scikit-learn where necessary. Random seeds were fixed for reproducibility. Across experiments:
  • Synthetic features $X_{1},X_{2}\sim\mathcal{N}(0,1)$ were generated independently for $n=1000$ observations unless otherwise stated.
  • Confounders $Z$ were constructed as linear combinations of $X_{1},X_{2}$ with Gaussian noise $\eta\sim\mathcal{N}(0,1)$.
  • Nonlinear transformations such as $\tanh(\cdot)$ and $\sin(\cdot)$ were used to define signal functions.
  • True outcomes $Y\sim\text{Bernoulli}(\sigma(f(X)+g(Z)))$ were generated given these transformations.
  • Logistic regressions were trained using only $X$, omitting $Z$, to simulate structural misspecification.
  • Portfolio weights were constructed as $\omega=2\hat{p}(X)-1$, where $\hat{p}$ is the predicted probability.
  • Optimization used empirical mean and covariance matrices to solve the classical Markowitz problem with quadratic programming (via CVXPY).
Monte Carlo experiments (Section 4.6) used n=200,000n=200{,}000 samples and varied $\sigma_{\zeta}\in[0,0.8]$. OLS regressions were computed with NumPy’s least squares solver. All empirical results were averaged over 10 repetitions.
A.1 Code for Calibration–Ranking Experiments
import os
import numpy as np
import matplotlib.pyplot as plt
np.random.seed(42)
os.makedirs("figures", exist_ok=True)
def make_spd_cov(n, n_factors=3, idio_scale=0.2):
    F = np.random.randn(n, n_factors)
    Lambda = np.diag(np.random.uniform(0.6, 1.4, size=n_factors))
    Sigma_f = F @ Lambda @ F.T
    D = np.diag(np.random.uniform(idio_scale, 2*idio_scale, size=n))
    V = Sigma_f + D
    V = 0.5*(V + V.T)
    w, Q = np.linalg.eigh(V)
    w = np.maximum(w, 1e-8)
    V = (Q * w) @ Q.T
    Vinv = (Q * (1.0/w)) @ Q.T
    return V, Vinv
def vminner(u, v, Vinv): return float(u.T @ Vinv @ v)
def vnorm(u, Vinv): return np.sqrt(max(vminner(u,u,Vinv), 1e-18))
def tangency_direction(signal, Vinv): return Vinv @ signal
def mean_vol_sharpe(w, mu, V):
    mean = float(mu.T @ w)
    vol = float(np.sqrt(max(w.T @ V @ w, 1e-18)))
    return mean, vol, mean/vol if vol>0 else np.nan
def generate_mu(n=100, scale=0.25):
    mu = np.random.randn(n)
    return mu / (np.linalg.norm(mu)+1e-12) * scale
def monotone_transforms(mu, powers):
    return [np.sign(mu) * (np.abs(mu)**p) for p in powers]
def main():
    n, powers = 120, (0.6, 0.8, 1.0, 1.2, 1.4)
    V, Vinv = make_spd_cov(n)
    mu = generate_mu(n=n, scale=0.25)
    tildes_raw = monotone_transforms(mu, powers)
    mu_norm = vnorm(mu, Vinv)
    tildes = [t*(mu_norm/(vnorm(t,Vinv)+1e-12)) for t in tildes_raw]
    # (a) Scatter
    plt.figure()
    for k in [0,2,4]:
        p = powers[k]
        plt.scatter(mu, tildes[k], s=12, alpha=0.55, label=f"p={p:.1f}")
    plt.axhline(0,color="black",lw=0.6,alpha=0.4)
    plt.axvline(0,color="black",lw=0.6,alpha=0.4)
[…]
    plt.title("Ranking preserved under monotone transforms")
    plt.legend(); plt.grid(True,ls="--",alpha=0.35)
    plt.tight_layout()
    plt.savefig("figures/ranking_preserved.png", dpi=220)
    # (b) Sharpe ratios
    rel_sharpes, sharpe_true = [], None
    for p, t in zip(powers,tildes):
        _,_,S_mu = mean_vol_sharpe(tangency_direction(mu,Vinv), mu, V)
        _,_,S_t  = mean_vol_sharpe(tangency_direction(t,Vinv), mu, V)
        if sharpe_true is None: sharpe_true = S_mu
        rel_sharpes.append(S_t/(S_mu+1e-12))
    plt.figure()
    plt.plot(powers, rel_sharpes, marker="o")
    plt.xlabel("Exponent \$p\$")
    plt.ylabel("Relative Sharpe")
    plt.title("Sharpe efficiency vs. calibration")
    plt.grid(True,ls="--",alpha=0.35)
    plt.tight_layout()
    plt.savefig("figures/sharpe_calibration.png", dpi=220)
if __name__=="__main__":
    main()
A.2 Simulated Experiment for Misspecified Signal Frontier
import numpy as np
import matplotlib.pyplot as plt
import cvxpy as cp
# Simulation parameters
np.random.seed(42)
n_assets = 5
n_obs = 1000
# Features and confounder
X = np.random.normal(0, 1, size=(n_obs, n_assets))
Z = 0.7 * X[:, 0] + 0.3 * X[:, 1] + np.random.normal(0, 1, size=n_obs)
logit_true = np.tanh(X) + 0.5 * np.sin(Z)[:, None]
p_true = 1 / (1 + np.exp(-logit_true))
# True returns
returns = (2 * p_true - 1) * 2.0 + np.random.normal(0, 0.05, (n_obs, n_assets))
# Misspecified logistic models per asset
betas = np.random.normal(1.0, 0.5, size=(n_assets, n_assets))
logits = X @ betas.T
p_pred = 1 / (1 + np.exp(-logits))
signals = 2 * p_pred - 1
# Empirical moments
mu_hat = signals.mean(axis=0)
Sigma = np.cov(returns, rowvar=False)
# Target returns and frontier
target_returns = np.linspace(mu_hat.min() * 1.5, mu_hat.max() * 1.5, 50)
efficient_risks = []
efficient_returns = []
for R in target_returns:
    w = cp.Variable(n_assets)
    prob = cp.Problem(cp.Minimize(cp.quad_form(w, Sigma)),
                      [cp.sum(w) == 1, mu_hat @ w == R])
    prob.solve()
    if w.value is not None:
        vol = np.sqrt(w.value @ Sigma @ w.value)
        efficient_returns.append(R)
        efficient_risks.append(vol)
# Plot
plt.figure(figsize=(8, 5))
plt.plot(efficient_risks, efficient_returns,
         label=’Misspecified Signals’, color=’tab:blue’)
plt.xlabel("Portfolio Volatility")
plt.ylabel("Expected Portfolio Return")
plt.title("Efficient Frontier from Misspecified Signals")
plt.grid(True)
plt.legend()
plt.tight_layout()
plt.savefig("efficient_frontier_qp_misspecified_scaled.png", dpi=300)
plt.show()
A.3 Efficient Frontier from Real Financial Data
import numpy as np
import pandas as pd
import matplotlib.pyplot as plt
import cvxpy as cp
# Load Excel file
file_path = "constituents_SPX_2008_Full.xlsx"
df = pd.read_excel(file_path)
# Date processing and returns
df[’Date’] = pd.to_datetime(df[’Date’], format=’%Y%m%d’)
df.set_index(’Date’, inplace=True)
returns = df.pct_change().replace([np.inf, -np.inf], np.nan).dropna()
# Subset selection: 5 assets over 1000 days
subset_returns = returns.iloc[:1000, :5]
mu_hat = subset_returns.mean().values
Sigma = subset_returns.cov().values
n_assets = len(mu_hat)
# Efficient frontier computation
target_returns = np.linspace(mu_hat.min() * 1.5, mu_hat.max() * 1.5, 50)
efficient_risks = []
efficient_means = []
for R in target_returns:
    w = cp.Variable(n_assets)
    constraints = [cp.sum(w) == 1, mu_hat @ w == R]
    problem = cp.Problem(cp.Minimize(cp.quad_form(w, Sigma)), constraints)
    problem.solve()
    if w.value is not None:
        port_return = mu_hat @ w.value
        port_vol = np.sqrt(w.value @ Sigma @ w.value)
        efficient_means.append(port_return)
        efficient_risks.append(port_vol)
# Plotting
plt.figure(figsize=(8, 5))
plt.plot(efficient_risks, efficient_means, label=’Efficient Frontier’, color=’tab:blue’)
plt.xlabel("Portfolio Volatility")
plt.ylabel("Expected Portfolio Return")
plt.title("Efficient Frontier (Real Data, Quadratic Optimization)")
plt.grid(True)
plt.legend()
plt.tight_layout()
plt.show()
A.4 Code for Bias Attenuation Simulation
import numpy as np
import matplotlib.pyplot as plt
# Parameters
alpha = 0.6
beta = 0.2
gamma = 0.7
n = 200_000
sigma_zeta_vals = np.linspace(0.0, 0.8, 17)
empirical_beta = []
theoretical_beta = []
# Loop over different confounding noise levels
for sigma_zeta in sigma_zeta_vals:
    X = np.random.normal(0, 1, size=n)
    eta = np.random.normal(0, np.sqrt(1 - alpha**2), size=n)
    zeta = np.random.normal(0, sigma_zeta, size=n)
    Z = -alpha * X + eta + zeta
    Z_tilde = Z / np.sqrt(1 + sigma_zeta**2)
    epsilon = np.random.normal(0, 1, size=n)
    Y = beta * X + gamma * Z_tilde + epsilon
    # OLS regression of Y on X alone
    X_centered = X - X.mean()
    beta_hat = np.dot(X_centered, Y - Y.mean()) / np.dot(X_centered, X_centered)
    empirical_beta.append(beta_hat)
    # Theoretical OVB formula
    beta_theory = beta - gamma * alpha / np.sqrt(1 + sigma_zeta**2)
    theoretical_beta.append(beta_theory)
# Plot results
plt.figure(figsize=(8, 5))
plt.plot(sigma_zeta_vals, empirical_beta, ’o’, label=’Monte Carlo’)
plt.plot(sigma_zeta_vals, theoretical_beta, ’--’, label=’Theory’)
[…]
plt.title("Bias Attenuation under Approximate Cancellation")
plt.legend()
plt.grid(True)
plt.tight_layout()
plt.savefig("appendix_E_bias_attenuation.png", dpi=300)
plt.show()
Appendix B Refutation of Structural Necessity Claims
Several recent critiques of predictive approaches to portfolio optimization (López de Prado et al.,, 2024, 2025) have advanced strong structural arguments suggesting that causal factor models are not only desirable but strictly necessary for efficiency. These arguments rest on claims such as: (i) misspecification inevitably produces sign inversion; (ii) sign inversion is the sole driver of inefficiency; (iii) sign agreement guarantees efficiency; and (iv) coherent efficient frontiers cannot exist without fully specified causal models.
The framework developed in this paper allows us to test these propositions formally. Across theory, controlled simulations, and empirical analysis, we show that these claims do not hold once more general conditions are considered. Misspecified models often preserve directional alignment, inefficiency frequently arises from calibration errors rather than inversion, and efficient frontiers remain convex provided that signals retain positive alignment with expected returns. Taken together, these results demonstrate that the essential determinants of viability are directionality and calibration, not structural fidelity. Misspecification can attenuate performance, but it does not by itself collapse optimization geometry or invalidate predictive frontiers.
Table 2 summarizes the principal claims, their logical form, and the corresponding refutations established in this paper. The recurring message is clear: predictive signals need not be causally identified to support coherent portfolio construction, and the mechanisms of inefficiency are more nuanced than previously asserted.
Claim from Prior Work Logical Form Refutation in This Paper
Misspecification necessarily causes sign inversion $\text{Misspecification}\Rightarrow\text{Sign Inversion}$ Counterexamples in Sections 3.2 and 3.4 show that alignment can persist under omitted-variable bias. Figures 1 and 3 confirm empirically that inversion is not inevitable.
Sign inversion is required for inefficiency $\text{Inefficiency}\Leftarrow\text{Sign Inversion}$ Theorem 1 proves that inefficiency arises from miscalibration even without inversion. Lemma 2 and Corollary 2 show that correct ranking with poor sizing remains inefficient.
Sign agreement guarantees efficiency $\text{Sign Agreement}\Rightarrow\text{Efficiency}$ Refuted by Theorem 1: aligned signals with distorted magnitudes yield attenuated Sharpe ratios. Efficiency requires both alignment and calibration.
Misspecification invalidates mean–variance optimization $\text{Misspecification}\Rightarrow\text{No Frontier}$ Lemma 4 and Corollary 3 prove that positive alignment suffices for smooth, convex frontiers. Section 4 (Figures 4, 5) confirms this empirically.
Only causal factor models support efficiency $\neg\text{Causal}\Rightarrow\neg\text{Efficient}$ Refuted theoretically and empirically: Section 4.5 shows that associational signals from lagged, non-causal features yield viable frontiers. Section 3 formalizes that efficiency depends on directional informativeness, not causal identification.
Misspecified signals produce anti-aligned weights $\omega^{\text{pred}}\approx-\omega^{\text{true}}$ No general basis. Figure 3 shows no clustering along the negative diagonal. Section 3.5 and Monte Carlo experiments (Section 4.6) demonstrate smooth attenuation, not inversion.
Table 2: Structural claims from (López de Prado et al.,, 2024, 2025) and their refutations. Across cases, inefficiency is shown to arise from miscalibration rather than inevitable inversion, and viable frontiers persist under positive alignment.
Appendix C Correction of Variance Normalization in Factor Models
In Section 3.1, we identified a normalization discrepancy in (López de Prado et al.,, 2024), where the reported variance of $F_{2}$ is inconsistent with standard factor modeling assumptions. We now provide the corrected derivation and its implications for bias interpretation.
C.1 Setup
Let:
$F_{2}=\delta Z+\eta,\quad Z\sim\mathcal{N}(0,1),\quad\eta\sim\mathcal{N}(0,1-\delta^{2}),\quad Z\perp\eta.$
Then $\mathrm{Var}(F_{2})=\delta^{2}+(1-\delta^{2})=1$, and $\mathrm{Cov}(Z,F_{2})=\delta$, maintaining unit variance and interpretable loadings.
C.2 Comparison
Assumption on $\mathrm{Var}(F_{2})$ Estimated Loading $\hat{\beta}_{n}$ Bias Characterization
Correct: $\mathrm{Var}(F_{2})=1$ $\hat{\beta}_{n}=\beta_{n}+\gamma_{n}\delta$ Linear in $\delta$; interpretable directional bias
(López de Prado et al.,, 2024): $\mathrm{Var}(F_{2})=1+\delta^{2}$ $\hat{\beta}_{n}=\dfrac{\beta_{n}+\gamma_{n}\delta}{1+\delta^{2}}$ Nonlinear attenuation; no longer proportionate to confounding strength
Table 3: Impact of variance normalization on estimated factor loadings and omitted-variable bias interpretation.
Despite functional differences, both formulations confirm that confounders induce signal bias. However, only the corrected form maintains scale-consistent estimation and analytically tractable bias propagation.
Appendix D Full Monte Carlo Results for Approximate Cancellation
This appendix complements Section 4.6 by reporting the full numerical table of simulation results for omitted-variable bias under increasing structural noise. See Appendix A for implementation.
$\sigma_{\zeta}$ $\hat{\beta}_{\mathrm{OLS}}$ (MC) $\beta_{\mathrm{theory}}$ Bias (MC) Bias (Theory)
0.000 -0.220295 -0.220000 -0.420295 -0.420000
0.050 -0.218724 -0.218690 -0.418724 -0.418690
0.100 -0.216333 -0.216330 -0.416333 -0.416330
0.150 -0.213370 -0.213374 -0.413370 -0.413374
0.200 -0.209934 -0.209951 -0.409934 -0.409951
0.250 -0.206124 -0.206150 -0.406124 -0.406150
0.300 -0.201989 -0.202021 -0.401989 -0.402021
0.350 -0.197564 -0.197600 -0.397564 -0.397600
0.400 -0.192882 -0.192920 -0.392882 -0.392920
0.450 -0.187965 -0.188005 -0.387965 -0.388005
0.500 -0.182833 -0.182874 -0.382833 -0.382874
0.550 -0.177504 -0.177544 -0.377504 -0.377544
0.600 -0.172000 -0.172039 -0.372000 -0.372039
0.650 -0.166337 -0.166373 -0.366337 -0.366373
0.700 -0.160533 -0.160566 -0.360533 -0.360566
0.750 -0.154606 -0.154636 -0.354606 -0.354636
0.800 -0.148573 -0.148600 -0.348573 -0.348600
Table 4: Monte Carlo and theoretical results for omitted-variable bias under approximate cancellation. Simulation size: n=200,000n=200{,}000. Parameters: $\alpha=0.6$, $\beta=0.2$, $\gamma=0.7$.
Statements and Declarations
Funding
The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.
Competing Interests
The authors have no relevant financial or non-financial interests to disclose.